OpenStudy (anonymous):

Evaluate the integral: sin^2(x)cos^3(x) dx

OpenStudy (anonymous):

Use \(u\) sub.

OpenStudy (anonymous):

okay so my u would be sin x right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Since you have \[ \int (\sin^2 x\cos^2x) \cos x\;dx = \int (\sin^2 x\cos^2x) \;d(\sin x) \]

OpenStudy (anonymous):

alright thanks i got it now